### v1

# Graphical Abstract Highlights

Abstract
West Nile virus (WNV) is a climate-sensitive mosquito-borne arbovirus circulating between mosquitoes of the genus Culex and birds, with a potential spillover to humans and other mammals. Recent trends in climatic change, characterized by early and/or prolonged summer seasons, increased temperatures, and above-average rainfall, probably facilitated the spread of WNV in Europe including Germany. In this work, we formulate a spatial WNV model consisting of a system of parabolic partial differential equations (PDEs), using the concept of diffusion and advection in combination with temperature-dependent parameters, i.e. mosquito biting rate, extrinsic incubation, and mortality rate. Diffusion represents the random movement of both mosquitoes and hosts across space, while advection captures the directed movement of migratory birds. The model is first studied mathematically, and we show that it has non-negative, unique, and bounded solutions in time and space. Numerical simulations of the PDE model are performed using temperature data for Germany (2019 - 2024). Results obtained from the simulation showed a high agreement with the reported WNV cases among birds and equids in Germany. The observed spreading patterns from the year 2018 to 2022 and the year 2024 were mainly driven by temperature in combination with diffusion processes of hosts and vectors. Only during the year 2023, the additional inclusion of advection for migratory birds was important to correctly predict new hotspots in new locations in Germany.
Keywords: West Nile virus, diffusion, advection, spread, mosquitoes, migratory birds, resident birds, hotspots
1 Introduction
West Nile virus (WNV) is an arbovirus of the Japanese encephalitis antigenic complex of the Flaviviridae family. It was first detected in a febrile patient from the West Nile district of Northern Uganda in 1937 (Smithburn et al., 1940), but is now considered the most geographically widespread arbovirus in the world (Ciota, 2017). The virus circulates in an enzootic cycle between Culex mosquitoes and birds, with regular spillovers to humans, equids, and other mammals (Calisher et al., 1989; Hubálek and Halouzka, 1999; Kramer et al., 2008). Birds are considered amplifying hosts as they develop sufficient levels of viremia, allowing for the infection of mosquitoes, while humans, equids, and other mammals are considered dead-end hosts (Rappole and Hubálek, 2003). Most humans infected with WNV are asymptomatic, but few may develop fever, headache, rash, among others (Petersen et al., 2013), including the risk of a fatal outcome (Campbell et al., 2002). Infected equids often exhibit ataxia, lethargy, and weakness as signs of the infection (Salazar et al., 2004). Multiple theories have been discussed regarding the introduction of new WNV cases in a previously unaffected region. There is evidence that migratory birds play an important role in the spread of the pathogen (Rappole and Hubálek, 2003). For example, birds captured during their seasonal migration have been tested positive for WNV in Israel (Ernek et al., 1977; Malkinson et al., 2002; Nir et al., 1967; Watson et al., 1972). Migratory birds likely spread the virus to stopover sites in Europe as they travel from the south to the north(Hannoun et al., 1972; Hubálek and Halouzka, 1999). In contrast, considering the speed and duration of flights of migratory birds in combination with the duration of viremia, other researchers argued that the short-distance movement of non-migratory birds better explains the observed spread pattern of WNV (Rappole and Hubálek, 2003).
Several mathematical models for WNV have been formulated. Many models emphasize the role of climatic conditions on the dynamics of WNV, as temperature plays a critical role in the life-history traits of mosquitoes and virus replication. Thereby, temperature-dependent functions are used to model mosquito and virus development rates (Heidecke et al., 2025; Laperriere et al., 2011; Mbaoma et al., 2024). These are included in ordinary differential equations (ODEs) to analyze the dynamics of WNV. However, local transmission dynamics are generally modeled within the vector and host populations without considering the geographic spread. Fesce et al. (2022) estimated the basic reproductive number of the WNV infection and used it to quantify the infection spread in mosquitoes and birds, to classify areas with temperatures suitable for the outbreak of WNV. Other studies focus on the control of WNV, e.g., Bowman et al. (2005) simulated mosquito reduction methods and personal protective measures. A potential spill-over from birds to equids and humans was modeled by Laperriere et al. (2011) using temperature-driven parameters and providing values for parameters that govern the developments and transitions from one compartment to another. Other models analyzed differences in seasonal trends using sinusoidal (trigonometric) functions to represent seasonal variations of the WNV disease dynamics (Abdelrazec et al., 2015; Cruz-Pacheco et al., 2009; Moschini et al., 2017). Such models usually show seasons when mosquitoes are most likely to be active.
Different modeling frameworks also include the role of migratory birds in the transmission of WNV. Bergsman et al. (2016) formulated an ODE model that investigates the interplay of WNV in both resident and migratory hosts. The model considers the yearly seasonal outbreaks that depend primarily on the number of susceptible migrant birds entering the local population each season, demonstrating that the early growth rates of seasonal outbreaks are more influenced by the migratory than the resident bird population. For Germany, just recently, Mbaoma et al. (2024) developed an ODE model for WNV in Germany by incorporating the complete life cycle of Cx. pipiens as a vector, resident birds, migratory birds, and humans as hosts. The study concluded that short-distance migratory birds such as the hooded crow were identified as important hosts in maintaining the enzootic transmission cycle of WNV in Germany.
Although ODE-based models allow us to understand the transmission dynamics of WNV, they do not generally consider the spatial movement of hosts and vectors across different regions without incorporating additional types of models, like multi-patch models or distance dispersal kernels. For example, Bhowmick et al. (2023) used an ODE-based modeling approach in combination with different distance dispersal kernels to describe the spreading process of WNV through migratory birds in Germany. As a promising alternative, researchers have incorporated the effects of spatial variations and disease spread across different geographical regions in ODE-based models, using partial differential equation (PDE) models. Lin and Zhu (2017) developed a reaction-diffusion PDE model for the spatial spread of WNV in mosquitoes and birds in North America using free-boundary conditions. Another study of the WNV propagation was performed by Maidana and Yang (2009), where they proposed a spatial model for analyzing the WNV spread across the USA, calculating the traveling wave speed. The model considers a system of PDEs that describe the movement of mosquito and avian populations using the diffusion and advection processes. Results from this study show that mosquito movements do not play a major role in the spread, while bird advection is an important factor for lower mosquito biting rates.
Maidana and Yang (2008, 2013) also contributed a lot to the spatial modeling of WNV using PDE models, and their framework is the basis of this study. Our model differs from previous WNV PDE models available in the literature, as we consider four different hosts, i.e., resident birds, migratory birds, humans, and equids, allowing us to capture the changes in the transmission patterns of WNV. Furthermore, we simulate the model in Germany, accounting for spatial heterogeneity, identify hotspots, and give insights into the contribution of migratory birds to WNV spread. In Germany, WNV was first detected in August 2018 in birds from the south-eastern region of Berlin (Ziegler et al., 2019). Since then, WNV has been established in Germany and has spread. Germany lies along migration routes of birds between Europe, the Mediterranean, and Africa (Musitelli et al., 2018). A study conducted by Nussbaumer et al. (2021) indicates that most migratory birds arrive in the northeastern region of Germany during spring and leave the country in early autumn via the southwestern direction. First WNV strains confirmed in Germany were closely related to strains previously detected in Austria and the Czech Republic (Ziegler et al., 2019), indicating a potential route of introduction, and their ability to overwinter in Cx. pipiens (Kampen et al., 2021) allows the long-term persistence of the virus.
2 Model formulation
Building on ODE-based models developed by Laperriere et al. (2011) and Mbaoma et al. (2024), we formulate a model for the spatial spread of WNV in Germany between 2019 and 2024. Our spatial model consists of a system of semi-linear parabolic PDEs governed by diffusion, advection, and reaction processes, together with temperature-dependent mosquito biting, latency, and mosquito mortality rates. In addition, our model includes the population of mosquitoes, resident and migratory birds. The population of mosquitoes at time and geographic location is divided into three compartments: susceptible exposed and infectious mosquitoes with the total population given by Susceptible mosquitoes get infected by biting an infectious resident bird and/or migratory bird with a temperature-dependent mosquito biting rate of
and a probability of acquiring the infection of and from resident and migratory birds respectively. The force of infection on mosquitoes is thus given by
where and are the environmental carrying capacities of the resident and migratory birds. A successful contact between infectious birds and susceptible mosquitoes at location and time sends susceptible mosquitoes to the exposed class. The latency rate on mosquitoes was fitted by Heidecke et al. (2025), and is given by
and from the exposed stage, mosquitoes progress to the infectious class Mosquitoes die at temperature-dependent natural mortality rate of where
defined by Laperriere et al. (2011). The host population consists of amplification hosts: resident birds and migratory birds. The population of hosts is divided into susceptible exposed , infected recovered and dead hosts, with and for easier readability, we denote resident birds by subscript and migratory birds by Laperriere et al. (2011) modeled the recruitment of susceptible birds using a gamma distribution, accounting for the observed seasonal birth rate of the American crow, which was the bird species observed to be mainly affected by the WNV in the Minneapolis metropolitan area. To reduce the non-linearity of our model, which has an impact on the numerical simulation of the PDE model, we compared our ODE simulation using a recruitment rate of but found no significant differences in the numerical results of our ODE version and theirs. Thus, our resident and migratory birds are recruited into the susceptible state at a rate of and respectively. Susceptible resident and migratory birds get infected by infectious mosquitoes, with a biting rate defined in Equation (1), and a probability of acquiring the infection of and respectively. The force of infection on resident birds is given by
while that of migratory birds is
The intrinsic incubation period of resident and migratory birds is given by and respectively, while their recovery rates are given by and respectively. The proportions of resident birds and of migratory birds recover, while the proportions and die at the rates and respectively. The natural death rate of resident birds and migratory birds is given by with
2.1 Diffusion and advection processes
Diffusion refers to the process by which matter is transported from one part of a system to another as a result of random molecular motions (Crank, 1979). The concept of diffusion dates back to the 19th century, when Fick (1855) recognized the transfer of heat by conduction due to random molecular motions and quantified this process by adopting the mathematical equation of heat conduction previously derived by baron de Fourier (1822). To date, the diffusion process has gained significant attention, with mathematical biology being one of the application areas. In ecological modeling, the diffusion process is used to capture the random movement of species from regions of higher concentration to regions of lower concentration (Murray, 2007), which has an impact on their spatial distribution patterns. On the other hand, the advection process refers to the directional flow of a substance by a known velocity vector field (Boyd, 2001). Diffusion and advection processes combined with the ODE terms form an advection-diffusion-reaction PDE system that models how a phenomenon spreads in both time and space. Reaction terms are responsible for the local transmission at each location, while diffusion and advection are responsible for the spatial movement of species, including infectious specimens. In this study, we define diffusion for mosquitoes and both resident and migratory birds, as they locally move in search of resting/breeding sites, food, and other needs. The advection process is only defined for migratory birds, as we can account for their seasonal directional migration in the northeast direction during spring to late spring and fly out in the southwest direction starting from autumn to late autumn (Nussbaumer et al., 2021). The diffusion term is denoted by the Laplacian operator with diffusion coefficients for mosquitoes, resident, and migratory birds denoted by in respectively. Moreover, we choose the diffusion coefficients in a way that the diffusion of mosquitoes is less than that of resident and migratory birds, as mosquitoes are considered insects with a small radius of action (Harrington et al., 2005; Maidana and Yang, 2009; Verdonschot and Besse-Lototskaya, 2014). The diffusion coefficient of resident birds should also be less than that of migratory birds, to highlight the higher mobility of migratory birds. The advection term is denoted by the gradient operator with a velocity deg/day, in a two-dimensional plane.
The schematic diagram of the model under study is shown in Figure (1), while the system of equations describing the dynamics of the WNV model under study is shown in system (7).
The total populations of mosquitoes, resident birds, migratory birds, humans, and equids are given by while the initial conditions are given by
where,
We assume the system moves in a region with a smooth boundary according to Fick’s law (Fick, 1855), so that in the initial conditions, where and subject to the homogeneous Neumann boundary conditions
and is the unit outer normal to
Based on available WNV data in Germany, birds of the Accipitridae family were the most affected bird species in 2018. Therefore, as also done by Mbaoma et al. (2024), we used parameters associated with northern goshawks and raptors for the resident bird, both of which fall under the Accipitridae family. Initial conditions of infectious residential birds are estimated to be high in Halle (Saale), the main WNV hot-spot in 2018, followed by Berlin, and lastly few birds in Laage, Poing, and Plessa. More precisely, there are 2 infectious resident birds in Berlin, 1 in Plessa, Poing, and Laage, and 7 in Halle (Saale), as described by the end-of-year 2018 cases (Ziegler et al., 2020). For simplicity, we assume the same locations for initial conditions of infectious migratory birds. In the initial condition of susceptible hosts and mosquitoes, we assume a spatially homogeneous distribution, and we assign the total populations of each host uniformly across the entire domain. The total initial population of susceptible resident birds is assumed to be 100,000,000, while susceptible migratory birds are set at 500,000. The total mosquito population is estimated to be 20 million susceptible mosquitoes and 80,000 infectious mosquitoes, which agrees with an approximate minimum infection rate of 0.355 per 1000 mosquitoes estimated by Bernard et al. (2001). For numerical stability, we set initial dead resident birds and migratory birds to 1.
The variables are summarized in Table (1).
Variable Definition Value susceptible mosquitoes exposed mosquitoes 0 infectious mosquitoes 80,000 susceptible resident birds exposed resident birds 0 infectious resident birds 12 recovered resident birds 0 dead resident birds 1 susceptible migratory birds 500,000 exposed migratory birds 0 infectious migratory birds 12 recovered migratory birds 0 dead migratory birds 1
The model parameters are summarized in Table (2).
Parameter Definition Value Source natural mortality rate of resident birds 0.00034 (Laperriere et al., 2011) natural mortality rate of migratory birds 0.00023 (Mbaoma et al., 2024) latency rate on mosquitoes Eqn (3) (Heidecke et al., 2025) latency rate on resident birds 0.196 (Mbaoma et al., 2024) latency rate on migratory birds 0.285 (Mbaoma et al., 2024) removal rate of infectious resident birds 0.867 (Mbaoma et al., 2024) removal rate of infectious migratory birds 0.4 (Mbaoma et al., 2024) proportion of dead resident birds 0.655 (Mbaoma et al., 2024) proportion of dead migratory birds 0.103 (Mbaoma et al., 2024) probability of a new resident bird infection 0.9 (Mbaoma et al., 2024) probability of a new migratory bird infection 0.97 (Mbaoma et al., 2024) probability of a new mosquito infection from a resident bird 0.4 (Mbaoma et al., 2024) probability of a new mosquito infection from a migratory bird 0.7 (Mbaoma et al., 2024) mosquito biting rate Eqn (1) (Rubel et al., 2008) natural mortality rate of mosquitoes Eqn (4) (Tran et al., 2013) carrying capacity of mosquito population defined by the authors carrying capacity of resident bird population defined by the authors carrying capacity of migratory bird population defined by the authors diffusion coefficient of mosquitoes Table (3) inferred from the data diffusion coefficient of resident birds Table (3) inferred from the data diffusion coefficient of migratory birds Table (3) inferred from the data advection vector for migratory birds section (6) inferred from the data
3 Mathematical analysis of the model
The mathematical properties of the model are shown in Appendix (8).
4 Steady states and the reproductive number
The PDE system (7) admits a non-trivial disease-free equilibrium point such that
The uniqueness of can be verified using [(Wang et al., 2018), Lemma 2.1]. The basic reproductive number of the deterministic part of our model is computed following van den Driessche and Watmough (2002), as follows:
The matrix of new infections is given by while that of transitions is given by where
Taking partial derivatives and evaluating at the disease-free equilibrium point, we have
and
The basic reproductive number is given by the spectral radius of which yields
where
and
Since all parameters of the model are positive, and the equilibrium point is also positive, then and are non-negative. From this, we observe that
with equality when which shows the contribution of migratory birds in increasing the amplifying host population. Figure (2) shows the plot of the basic reproductive number all over Germany, as of 2024. shows a hotspot in the south-eastern region of Germany towards the northern part, which is in agreement with the observed data. It also identifies another hotspot in the south-west, around Bobenheim-Roxheim. The regions where is zero are regions where the temperature data has no values, mainly because of the presence of a sea. To validate the simulated we performed a chi-squared goodness of fit test at the 5% significance level. From this test, we obtained a p-value of which indicates that there are no significant differences between the observed and simulated patterns.
5 Simulation framework
The PDE model is simulated starting from 1 January 2019 until 5 December 2024 using a daily time step. Cases for the year 2018 (Figure 2(a)) are set as initial conditions at each location where cases were confirmed by 1 January 2019. To account for the inflow of new data through ongoing WNV surveillance activities, we model the spread every year, in such a way that the initial conditions for each year are the final solution of the previous year. This allows us to update the initial conditions in areas where the model did not predict cases, e.g. in 2019 a new case of WNV was confirmed in the north of Germany (Hamburg), and cases in this region may have played a role in the spreading pattern observed in the north of Germany during the subsequent years. In the process, we use inverse modeling to infer the diffusion and advection coefficients that could have contributed to the spreading pattern for each year. Temperature strongly influences the activity of Cx. pipiens, including their competence in transmitting WNV (Ciota et al., 2014), and thus, we included a temperature-dependent mosquito biting rate latency rate and mortality rate Hourly 2m temperature data from the ERA5 reanalysis is downloaded from the climate data store (Hersbach et al., 2023) and averaged to yearly data. During the simulation process, and are dynamically updated based on the temperature at different locations in Germany. To solve our PDE model, we use the Matlab PDEToolbox (The MathWorks, Inc., 2024), which uses a finite element algorithm. The toolbox solves PDEs of the form
and in our case, and matrices and are defined in Appendix (8.2). The matrix consists of the diffusion coefficients, while stores advection terms. The model is solved as follows (The MathWorks, Inc., 2024):
The map of Germany is downloaded as a shape file from (https://gadm.org/download_country_v3.html) and simplified using mapshaper (Harrower and Bloch, 2006). WNV infection data consisting of infected and dead birds and equids for the years 2018 - 2024 (Figure (2(a))) were obtained from the World Organisation for Animal Health (2018 - 2024) (WOAH). Human WNV cases data was obtained from European Centre for Disease Prevention and Control (2019, 2020, 2021, 2022, 2023, 2024). The observed data indicates presence data only, i.e., no absence data is available. For this reason, we qualitatively present our results (Low - High) to visualize the simulation results. To assess how well the model predictions match the observational data, we performed a chi-squared goodness of fit test at 5% level of significance, with the hypothesis:
There are no significant differences between the observed and simulated patterns.
There are significant differences between the observed and simulated patterns.
p-values are computed and the distributive patterns are tested for each year, while the null hypothesis is rejected whenever p-values are less than or equal to 0.05. The relationship between the model output and the observed data is further quantified using the cosine similarity index which measures the extent to which two vectors point in the same direction, without comparing their magnitudes (Huang et al., 2008). This index is computed using the formula where and represent the two datasets (modeled and observed) expressed as vectors. It is measured in the interval , where values closer to 1 indicate a high degree of similarity between the two vectors, while values close to zero indicate no similarity in the patterns described by these vectors. Values closer to -1 indicate that the two vectors point in opposite directions (Huang et al., 2008). Initial conditions and parameter values used in the simulations are summarized in Tables (1) and (2), respectively. We first simulate the model with migratory birds. It is indisputable that at every location with infectious resident and migratory birds or equids, infectious mosquitoes are also present. Therefore, we validated the model prediction using the distribution pattern of infectious resident birds
6 Results
The model initialized with 2018 WNV cases data shows strong visual agreement with observed patterns in Germany throughout the 2019 - 2024 simulation period, especially predicting hotspots in the eastern regions, including Berlin. The chi-squared test for 2019 indicated significant differences (p = 0.01), likely due to the model missing new cases in northern and central Germany; the cosine similarity was still moderately high at 0.6803 (Table 3). In subsequent years (2020 - 2024), p-values exceeded 0.05. Cosine similarity improved over the years, reaching 0.99 in 2024. Diffusion coefficients varied annually, showing a slight increase over the years, indicating higher mobility of both mosquito and host populations. Notably, in 2023 and 2024, the model accurately predicted cases in regions where none had been reported previously, hinting at its ability to anticipate under-reported spread patterns (Figure 4(e) - 4(f)). This capability was supported when 2024 observations confirmed some of the simulated 2023 circulation areas. The model’s performance suggests that incorporating spatial heterogeneity significantly enhances its predictive power. The diffusion speeds of both mosquitoes and hosts were positively correlated to temperature (Figure 5).
Year - squared statistical test Cosine similarity index Diffusion coefficient Value 0.0001 2019 <0.05 0.75 0.0003 0.0005 0.0002 2020 >0.05 0.68 0.00035 0.0004 0.00045 2021 >0.05 0.77 0.0005 0.00056 0.0005 2022 >0.05 0.73 0.00065 0.0007 0.05 2023 >0.05 0.80 0.06 0.07 0.05 2024 >0.05 0.99 0.065 0.07
6.1 Impact of migratory birds in disseminating WNV to distant regions:
During the years 2019 - 2024, the model is initially simulated with migratory birds. The advection vector is set to zero during the winter and summer seasons, as we assume that no migration is taking place. We then set the direction of flight to the northeast direction during spring, i.e., 1 March to 31 May of each year. During the autumn (1 September to 30 November), migratory birds start migrating towards the southwest, and we set the vector to indicate a southwest flow each year. Model results were the same for the cases when there were no migratory birds for all the years, except for the year 2023. During the year 2023, without migratory birds (Figure 6(a)), the model could not account for the new isolated cases observed in the western parts of Germany. After introducing migratory birds and their seasonal flow (Figure 6(b)), the model correctly predicts the observed pattern of WNV in 2023.
7 Discussions and conclusions
Understanding the WNV spreading pattern allows early detection of WNV outbreaks. The identification of hotspots also helps early planning and resource allocation for preventive measures. At the same time, it is expensive to carry out surveillance programs aiming to understand the spread of WNV. Therefore, in this study, we modeled the spread of WNV in Germany using a temperature-driven PDE model with advection and diffusion terms. In agreement with previous studies (Bhowmick et al., 2020; Bergsman et al., 2016; Mbaoma et al., 2024), at the temporal level, the basic reproductive number indicates that the presence of migratory birds increases the number of amplifying local hosts. At spatial scale, identified low - high - risk regions of public health interest, which agreed with the observed WNV cases in birds and equids. The reproductive number classified regions in the eastern parts of Germany as the main hotspot, while over time, new circulation areas were identified in the north and south. In addition, this study draws insights into the spreading pattern of WNV in Germany. Our approach, which uses the diffusion process of mosquitoes and birds, and advection for migratory birds in combination with temperature-dependent parameters, can infer the spreading speed of WNV in Germany. The simulation of the PDE model shows that the diffusion coefficients are high during the years 2023 and 2024. Due to lack of real movement data for hosts, we chose the diffusion coefficients that made the model prediction match the observed data. During the year 2023, the model prediction showed a poor fit to observed data when only resident birds were considered. However, incorporating migratory bird compartments, and capturing their directional seasonal flyways led to an improved fit for the same year. Because the pattern of observed WNV cases aligns with known migratory bird flyways, this result suggests that directed long-range bird migration could have played a significant role in transporting WNV across different locations in Germany, as also underlined by a recent study by Tóth et al. (2025). Our results imply that migratory birds may still have played an indirect role in shaping the WNV spreading pattern in Germany, even if they may not be the primary source of infection. Thus, effective surveillance and prevention efforts of WNV should also be implemented at key stopover sites, where migratory birds are likely to rest and interact with local mosquitoes. The observed data for 2023 did not validate regions in the southwest of Germany as potential hotspots, as only one single equid was confirmed in 2023 in this region. Interestingly, in 2024, the data obtained confirmed the southwestern part as a WNV circulation area, highlighting the predictive capacity of our model to accurately predict the WNV spreading pattern in Germany. The current WNV surveillance system remains patchy, compared to active surveillance (entomological, animal sentinels or blood donations) and passive surveillance systems (dead birds or symptomatic humans) that are cheaper and simpler, but are unable to detect the virus early before human cases occur. This is evidenced by that most data arises from incidental or passive detection rather than systematic monitoring (Constant et al., 2022).
Results from our spatially explicit PDE model demonstrate that temperature has a significant influence on the spreading pattern of WNV. Despite our model’s ability to simulate the spread of WNV in Germany, some additional factors can be considered to improve such PDE models in the future. For example, land use is known to influence the general distribution of mosquito populations (Watts et al., 2021), which plays a huge role in the mosquito abundance. Furthermore, we used temperature only to explain the yearly mosquito biting and mortality rates. Precipitation, wind speed, and direction also influence the general distribution and activity of mosquitoes (Ducheyne et al., 2007). The estimation of diffusion and advection coefficients by inverse modeling causes uncertainties in the results of the model, and thus, this work can be improved by considering more systematic methods for inferring diffusion and advection coefficients, such as Bayesian methods (Barajas-Solano and Tartakovsky, 2019). To further capture the random movement of hosts and mosquitoes, this PDE model can be extended to a stochastic PDE system to study more random effects that cannot be modeled by deterministic PDEs.
CRediT authorship contribution statement
Pride Duve: Designed the study, developed the model, proved the mathematical properties, computed the numerical simulations, and wrote the original draft. Renke Lühken: Supervised the study, designed the model, and reviewed the final draft. Felix Sauer: Supervised the study, designed the model, and reviewed the final draft.
Declaration of competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The cleaned WNV data for birds and equids used in this manuscript is freely available Here. Original datasets can be obtained from the World Organisation for Animal Health (2018 - 2024) (WOAH) website. The original data for human WNV cases is available from the ECDC website (European Centre for Disease Prevention and Control, 2019, 2020, 2021, 2022, 2023, 2024).
Acknowledgements
The authors would like to thank the Federal Ministry of Education and Research of Germany (BMBF) under the project NEED (Grant Number 01Kl2022), the Federal Ministry for the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (Grant Number 3721484020), and the German Research Foundation (JO 1276/51) for funding this project. The authors are also grateful to the National Research Platform for Zoonosen for funding the research visit of Pride Duve to Prof. Dr. Reinhard Racke’s lab, at the University of Konstanz. Furthermore, we would like to acknowledge and thank Reinhard Racke, Nicolas Schlosser, and Stefanie Frei for the discussions, consultations, and support.
8 Appendix
Theorem 1.
System (7) has non-negative and unique solutions that are bounded in
Proof.
We write system (7) together with the corresponding initial conditions, in the Banach space of continuous functions that include the boundary, as follows (Abboubakar et al., 2023):
where
and
Next, we show that is locally Lipschitz in Thus we show that
where is the Lipschitz constant. We observe that
Simplifying, we obtain
By the triangular inequality, and some simplifications, we arrive at
and thus is locally Lipschitz in By [(Capasso, 1993), Theorem B.17], (Mora, 1983), and [(Smoller, 1983), Theorem 14.4], there exist a local smooth and unique solution of system (7) in We observe that system (7) can be written in the form of system (14.12) in the book (Smoller, 1983), together with initial data defined in system (14.13). By [(Smoller, 1983), Theorem (14.14)], the solutions of system (7) are always positive. ∎
8.1 Well-posedness of the model in a feasible region
Theorem 2.
System (7) is well-posed mathematically and biologically, and for any
system (7) admits a unique positive solution:
satisfying:
Moreover, there exist another constant independent of initial data, such that the solution satisfies:
Proof.
We follow the approach presented in (Peng and Zhao, 2012; Wang and Dai, 2022). Using the regularity theory of parabolic PDEs (Pao, 1993), system (7) admits a unique non-negative classical solution
where represents the maximal existence time of the solution. By the strong maximum principle (Protter and Weinberger, 1984), then where are positive in
Summing up the equations and integrating over
we get
and are bounded for all By the positivity of and [(Peng and Zhao, 2012), Lemma (3.1)], with we conclude that there exist a positive constant that does not depend on initial data such that the solution satisfies
Similar arguments can be made for the mosquito population, thus we conclude that system (7) is well-posed. ∎
8.2 Matrices of the PDEToolbox
while